Find the point-slope form of an equation of a line through the points (-1,4) and (-3/5, 3/2)
I am not getting the this problem, and the answer in the book does not make sense. Please Help

Answer by lwsshak3(11628)

Find the point-slope form of an equation of a line through the points (-1,4) and (-3/5, 3/2)
Point-slope formula: y1-y2=m(x1-x2)
m=∆y/∆x=(4-3/2)/(-1-(-3.5))=(5/2)/(-2/2)=-25/4
using one of the points (-1,4)
y+1=-(25/4)(x-4)